An electron experiences an electric force of 1.8x10-11N at a distance of 5.0x10-9 m from the nucleus of an ion. The electron is moved farther away, to a distance of 2.0x10-8 m from the ion. What is the new electric force on the electron?




1
Expert's answer
2021-02-22T10:24:41-0500
"F=k\frac{qQ}{r^2}\\F'=k\frac{qQ}{r'^2}=F\frac{r^2}{r'^2}\\F'=1.8\cdot10^{-11}\frac{5^2}{20^2}=1.125\cdot10^{-12}\ N"